[Hallc_running] [New Logentry] Hall Status and plan

ciprian at jlab.org ciprian at jlab.org
Sun Oct 10 09:15:01 EDT 2021

Logentry Text:
Up to ~2am this morning (https://logbooks.jlab.org/entry/3921021) an IOC used by the PSS system went down and that dropped the hall to restricted. To remedy this MCC called in RadCon to do a sweep, which was completed at around 4am (https://logbooks.jlab.org/entry/3921056). Afterwards due to various issues beam delivery to hall C was impeded. MCC now believes all of those things are solved and have delivered 20uA to the carbon target. The beam was in the correct location so we moved to the production target and restarting the data taking.

The plan is to continue straight on until tomorrow morning at 9AM when Hall A setup is supposed to begin. We will meet tomorrow at 10AM for the RC daily discussion.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3921252
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211010/d4e496f2/attachment.html>

More information about the Hallc_running mailing list